A generator produces 36.0 V when turning at 940 rev/min. What emf does it produce when turning at 550 rev/min?
Answer #1
the generator produced votage should be propotional to fequency if the remining parameters are the same.
e (voltage ) propotional to n (frequency)

thus, e1 /e2 = n1/n2

36/e2 = 940/550

e2 =36(550)/940

= 21.06 V